OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the first equation is substituted into the second equation. 12x – y = –4 4x – 3y = –6

OpenStudy (anonymous):

from the first equation, 12x - y = -4 we take -y to the other side 12x = -4 + y (you get the sign change right?) then we take the -4 to the other side 12x + 4 = y (1) Now put value of y from (1) in the other equation => 4x - 3(12x +4) = -6 4x - 36x -12 = -6 -32x = -6 + 12 -32 x = -6 x = 6/32 = 3/16